### v2

# Injective envelopes of $C^{*}$-algebras as maximal rigid multiplier covers

Abstract.
We present a noncommutative analogue of Błaszczyk’s elegant two-step construction of the Gleason cover: first maximise irreducibility, then compactify. The key insight is that extremal disconnectedness (respectively, injectivity) emerges naturally from a maximality principle rather than being imposed by hand.
We work with $A$-multiplier covers: non-unital $C^{*}$-algebras $E$ equipped with a faithful, nondegenerate *-homomorphism $\iota:A\to M(E)$. Covers are ordered by $A$-preserving $\mathrm{ucp}$ maps between multiplier algebras in the forward direction. A cover is rigid if every $\mathrm{ucp}$ endomorphism of $M(E)$ fixing $A$ pointwise is the identity. We prove that if $(E_{\max},\iota_{\max})$ is maximal among rigid covers, then the inclusion $A\hookrightarrow M(E_{\max})$ is rigid and essential; consequently,
$I(A)\ \cong\ M(E_{\max})$
canonically over $A$ by Hamana’s characterisation. In the commutative case $A=C(X)$, we recover Błaszczyk’s picture: for a maximal regular refinement $(X,T^{\ast})$ with irreducible identity, $G(X)=\beta(X,T^{\ast})$ and $I(C(X))\cong C(G(X))\cong M(C_{0}(U))$ for any dense open $U\subseteq G(X)$. Thus the paradigm maximise first, then compactify provides a unified conceptual framework on both sides of Gelfand duality.
Key words and phrases:
Injective envelope, multiplier algebra, extremally disconnected space, Gleason cover, rigid extension
2020 Mathematics Subject Classification:
Primary 46L05; Secondary 54G05, 46L30
Institute of Mathematics, Czech Academy of Sciences (RVO: 67985840). Funding received from NCN Sonata-Bis 13 (2023/50/E/ST1/00067) is acknowledged with thanks.
1. Introduction
The injective envelope of a $C^{*}$-algebra, introduced by Hamana in his foundational papers [5, 6], has become a central tool in noncommutative analysis. For a compact Hausdorff $X$, injectivity on the commutative side amounts to extremal disconnectedness: $I(C(X))\cong C(K)$ with $K$ extremally disconnected. Gleason’s classical construction [4] of $K$ is intricate; Błaszczyk [1] later gave a strikingly concise route: first maximise the regular topology under an irreducibility constraint, then compactify to obtain $G(X)$. Indeed, in Błaszczyk argument, maximality is the driver and extremal disconnectedness is the consequence.
Our aim is to transpose this to the noncommutative setting. The multiplier algebra $M(E)$ is the natural analogue of the Čech–Stone compactification: $M(C_{0}(Y))\cong C(\beta Y)$. We introduce $A$-multiplier covers $(E,\iota)$ and a rigidity notion paralleling Hamana’s. The punchline is that a maximal rigid cover forces $M(E)$ to be a rigid essential extension of $A$, hence identifies canonically with $I(A)$.
Dictionary.
Topology $C^{\ast}$-algebra
irreducible surjection $Y\twoheadrightarrow X$ rigid/essential embedding $A\hookrightarrow B$
extremally disconnected compact $K$ injective $C^{\ast}$-algebra $B$
Gleason cover $G(X)$ injective envelope $I(C(X))$
Čech–Stone compactification $Y\mapsto\beta Y$ multiplier algebra $E\mapsto M(E)$
maximal regular refinement $T^{\ast}$ maximal rigid $A$-multiplier cover $(E_{\max},\iota_{\max})$
[…]
Organisation. Section 2 recalls Hamana’s characterisation of $I(A)$ and the multiplier-ideal technology we need. Section 3 introduces multiplier covers, defines the forward order by AA‑preserving $\mathrm{ucp}$ maps, and proves that maximality forces essentiality via a lifting argument through $I(A)$. It also contains an intrinsic transfinite existence proof of maximal rigid covers using minimal $A^{1}$‑projections and the Choi-Effros product. The main identification $M(E_{\max})\cong I(A)$ is established in Section 4, and Section 5 revisits the commutative picture, recovering Błaszczyk. We end with remarks on functoriality and local multipliers.
2. Preliminaries
2.1. Hamana’s injective envelope
An operator system $\mathsf{E}$ is injective if any $\mathrm{ucp}$ map $\phi:\mathcal{S}\to\mathsf{E}$ from an operator system $\mathcal{S}\subseteq\mathcal{T}$ extends to a $\mathrm{ucp}$ map $\tilde{\phi}:\mathcal{T}\to\mathsf{E}$.
Theorem 2.1.
For every operator system $A$ there exists an injective operator system $I(A)$ and a unital complete order embedding $j:A\hookrightarrow I(A)$ such that:
  1. (a)
    if $A\subseteq\mathsf{E}\subseteq I(A)$ with $\mathsf{E}$ injective, then $\mathsf{E}=I(A)$;
  2. (b)
    if $\Phi:I(A)\to I(A)$ is $\mathrm{ucp}$ and $\Phi|_{A}=\mathrm{id}$, then $\Phi=\mathrm{id}$.
Moreover, $I(A)$ is unique up to a unique complete order isomorphism over $A$. If $A$ is a $C^{\ast}$-algebra, then $I(A)$ is a $C^{\ast}$-algebra and $j$ is a *-monomorphism.
Equivalently, $I(A)$ is the unique rigid essential extension of $A$: whenever $q:I(A)\to\mathsf{F}$ is a complete quotient map and $q|_{A}$ is a complete isometry, then $q$ is injective.
2.2. Multiplier algebras and quotienting ideals
If $E$ is a (not necessarily unital) $C^{\ast}$-algebra, its multiplier algebra $M(E)$ is the $C^{\ast}$-algebra of double centralisers; see Busby [2] or Pedersen [8, Ch. 3]. If $Y$ is locally compact Hausdorff, then $M(C_{0}(Y))\cong C_{b}(Y)\cong C(\beta Y)$.
We shall use the following standard fact.
Lemma 2.2.
Let $E$ be a $C^{\ast}$-algebra and $I\lhd M(E)$ a closed two-sided ideal. Put
[…]
Then the quotient map $q_{K}:E\to E/K$ induces a surjective *-homomorphism
$\pi_{K}:\ M(E)\longrightarrow M(E/K)$
whose kernel contains $I$. Conversely, for a closed ideal $K\lhd E$, the strict extension $M(E)\twoheadrightarrow M(E/K)$ has kernel $\{m\in M(E):mE\subseteq K,\ Em\subseteq K\}$.
Proof.
See [8, §3.12] and [2]. For $x\in I$ one has $xE,Ex\subseteq K$, hence $x\in\ker\pi_{K}$. The converse kernel description is the usual characterisation of multipliers vanishing modulo $K$. ∎
Remark 2.3.
If $B$ is unital and $J\lhd B$, the canonical map $B\to M(J)$ given by left/right multiplication is a surjective *-homomorphism with kernel $\mathrm{Ann}(J):=\{b\in B:bJ=Jb=0\}$; hence
$M(J)\ \cong\ B/\mathrm{Ann}(J).$
Thus $M(J)=B$ iff $J$ is essential in $B$. In particular, even if $B$ is a $W^{*}$-algebra (e.g. a product of factors), $M(J)$ need not be $W^{*}$, because $\mathrm{Ann}(J)$ need not be weak-closed. This is consistent with Hamana’s examples where $I(A)$ is monotone complete but not von Neumann [7].
3. Multiplier covers and maximality
Definition 3.1.
Let $A$ be a $C^{*}$-algebra. An $A$-multiplier cover is a pair $(E,\iota)$ where $E$ is a (not necessarily unital) $C^{*}$-algebra and $\iota:A\to M(E)$ is a faithful, nondegenerate *-homomorphism; we identify $A$ with $\iota(A)\subseteq M(E)$. The cover is rigid if every $\mathrm{ucp}$ map $\Phi:M(E)\to M(E)$ with $\Phi|_{A}=\mathrm{id}_{A}$ equals the identity.
Given two $A$-multiplier covers $(E_{1},\iota_{1})$ and $(E_{2},\iota_{2})$, we write
$(E_{1},\iota_{1})\ \preccurlyeq\ (E_{2},\iota_{2})$
[…]
Intrinsic existence via minimal $A^{1}$-projections
Throughout this subsection we embed $A$ faithfully and nondegenerately into $\mathcal{B}(H)$ and write $A^{1}$ for its unitisation inside $\mathcal{B}(H)$. By an $A^{1}$-map we mean a unital completely positive map fixing $A^{1}$ pointwise. An $A^{1}$-projection is an idempotent $A^{1}$-map. We endow the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology. In what follows, the only continuity we use is:
$\Phi_{\alpha}\to\Phi\ \ \text{point-ultraweak}\quad\Longrightarrow\quad\Phi_{\alpha}\circ T\ \to\ \Phi\circ T\ \ \text{point-ultraweak}$
for each bounded linear $T:\mathcal{B}(H)\to\mathcal{B}(H)$; this is automatic because $(\Phi\mapsto\Phi\circ T)$ is continuous when evaluated on fixed vectors.
Lemma 3.2.
[…]
$P_{\lambda}\circ P_{\lambda_{0}}\;=\;P_{\lambda}\;=\;P_{\lambda_{0}}\circ P_{\lambda}.$
Then there exists a ucp map $P_{\infty}:\mathcal{B}(H)\to\mathcal{B}(H)$ with $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$ such that, for every $\lambda_{0}\in\Lambda$,
$P_{\infty}\circ P_{\lambda_{0}}\;=\;P_{\infty}\;=\;P_{\lambda_{0}}\circ P_{\infty}.$
[…]
Proof.
Equip the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology. By compactness of the unit ball of completely bounded maps, a subnet of $(P_{\lambda})_{\lambda}$ converges point-ultraweakly to a ucp map $P_{\infty}$; pointwise convergence preserves $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$. Fix $\lambda_{0}$ and $x\in\mathcal{B}(H)$. Using $P_{\lambda}\circ P_{\lambda_{0}}=P_{\lambda}$ for all $\lambda\geq\lambda_{0}$,
$P_{\infty}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}(x)\;=\;P_{\infty}(x),$
so $P_{\infty}\circ P_{\lambda_{0}}=P_{\infty}$. A symmetric limit using $P_{\lambda_{0}}\circ P_{\lambda}=P_{\lambda}$ gives $P_{\lambda_{0}}\circ P_{\infty}=P_{\infty}$. Hence $P_{\infty}$ is a lower bound. ∎
Proposition 3.3.
There exists a minimal $A^{1}$–projection $P:\mathcal{B}(H)\to\mathcal{B}(H)$ for the order
$\Phi\preccurlyeq\Psi\quad\Longleftrightarrow\quad\Phi\circ\Psi=\Phi=\Psi\circ\Phi\quad(\text{equivalently, }\operatorname{Ran}\Phi\subseteq\operatorname{Ran}\Psi).$
Moreover, any such minimal $P$ is idempotent: $P^{2}=P$.
Proof.
[…]
Lemma 3.4.
Let $P$ be minimal and let $\Xi$ be any $A^{1}$-projection. Then $P\circ\Xi\circ P=P$.
Proof.
Put $F:=P\circ\Xi\circ P$. Then $F$ is an $A^{1}$-projection and $F\preccurlyeq P$, so by minimality $F=P$. ∎
Lemma 3.5.
Let $P$ be minimal and $R:=P(\mathcal{B}(H))$ (an operator system). If $Q:R\to R$ is a ucp idempotent with $Q|_{A^{1}}=\mathrm{id}$, then $Q=\mathrm{id}_{R}$.
Proof.
Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be the inclusion and set $\widetilde{Q}:=\kappa\circ Q\circ P$. Then $\widetilde{Q}$ is an $A^{1}$-projection, so by Lemma 3.4,
$P\ =\ P\circ\widetilde{Q}\circ P\ =\ P\circ\kappa\circ Q\circ P\ =\ Q\circ P,$
using $P\circ\kappa=\mathrm{id}_{R}$ and $P^{2}=P$. For $r=P(x)\in R$ we get $Q(r)=Q(P(x))=P(x)=r$. ∎
Proposition 3.6.
With $P$ and $R=P(\mathcal{B}(H))$ as above, every ucp map $\theta:R\to R$ with $\theta|_{A^{1}}=\mathrm{id}$ equals $\mathrm{id}_{R}$.
Proof.
Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be inclusion and set $\Xi:=\kappa\circ\theta\circ P$. Then $\Xi$ is an $A^{1}$-projection and by Lemma 3.4,
$P\ =\ P\circ\Xi\circ P\ =\ P\circ\kappa\circ\theta\circ P\ =\ \theta\circ P.$
Thus for $r=P(x)\in R$ we have $\theta(r)=\theta(P(x))=P(x)=r$. ∎
Lemma 3.7.
If $C$ is a $C^{\ast}$-algebra and $P:C\to C$ is a ucp idempotent, then $\operatorname{Ran}P$ becomes a $C^{\ast}$-algebra with product $x\odot y:=P(xy)$ and the inherited involution, and the inclusion $\operatorname{Ran}P\hookrightarrow C$ is a complete order isomorphism.
Proof.
This is the Choi-Effros theorem [3, Thm. 3.1]. ∎
Proposition 3.8.
Let $P$ be minimal and equip $E:=P(\mathcal{B}(H))$ with the Choi-Effros product. Then:
  1. (a)
    $E$ is a $C^{\ast}$-algebra and $A$ embeds as a $C^{\ast}$-subalgebra (the product on $A$ agrees with the original one).
  2. (b)
    The inclusion $A\hookrightarrow E$ is rigid.
  3. (c)
    The operator system $E$ is injective.
  4. (d)
    The inclusion $A\hookrightarrow E$ is essential in the quotient sense: if $q:E\to F$ is a complete quotient and $q|_{A}$ is a complete isometry, then $q$ is injective.
Proof.
(a) For $a,b\in A^{1}$, $a\odot b=P(ab)=ab$ as $P|_{A^{1}}=\mathrm{id}$. In particular, the restriction to $A$ is a *-homomorphism.
(b) This is Proposition 3.6 applied to $R=E$.
(c) Let $\mathcal{S}\subseteq\mathcal{T}$ be operator systems and $u:\mathcal{S}\to E$ ucp. Compose $u$ with the inclusion $E\hookrightarrow\mathcal{B}(H)$ and extend to $\tilde{u}:\mathcal{T}\to\mathcal{B}(H)$ by injectivity of $\mathcal{B}(H)$. Then $P\circ\tilde{u}:\mathcal{T}\to E$ is a ucp extension of $u$.
(d) Let $q:E\to F$ be a complete quotient with $q|_{A}$ completely isometric. Let $\sigma_{0}:q(A)\to E$ be the composition $(q|_{A})^{-1}:q(A)\to A$ followed by the inclusion $A\hookrightarrow E$; then $\sigma_{0}$ is ucp. By (c), $\sigma_{0}$ extends to a ucp map $\sigma:F\to E$ with $\sigma\circ q|_{A}=\mathrm{id}_{A}$. Hence $\sigma\circ q:E\to E$ is ucp and fixes $A$, so by (b) it equals $\mathrm{id}_{E}$. Thus $q$ is injective. ∎
Theorem 3.9.
For every $C^{\ast}$-algebra $A$ there exists a maximal rigid $A$-multiplier cover in the order of Definition 3.1.
Proof.
Construct $E:=P(\mathcal{B}(H))$ via a minimal $A^{1}$-projection and endow it with the Choi-Effros product as before. By Proposition 3.8, $A\subseteq E$ is rigid and essential, and $E$ is injective as an operator system. Let $J\triangleleft E$ be an essential ideal (take $J=E$ if $E$ is unital). Then $M(J)\cong E$ (Remark 2.3), and the inclusion $\iota:A\hookrightarrow M(J)$ yields a rigid $A$-multiplier cover $(J,\iota)$.
We show $(J,\iota)$ is maximal. Let $(F,\jmath)$ be any rigid $A$-multiplier cover with
$(J,\iota)\ \preccurlyeq\ (F,\jmath),$
so there exists an $A$-preserving $\mathrm{ucp}$ map $\Theta:M(J)\to M(F)$ with $\Theta\circ\iota=\jmath$. Since $M(J)\cong E$ is injective as an operator system, there is an $A$-preserving $\mathrm{ucp}$ map $\Sigma:M(F)\to M(J)$ with $\Sigma\circ\jmath=\iota$ (Arveson extension along $\jmath$ into $E$).
Now
$(\Sigma\circ\Theta)\circ\iota\;=\;\Sigma\circ\jmath\;=\;\iota,$
hence $\Sigma\circ\Theta$ is a $\mathrm{ucp}$ endomorphism of $M(J)$ fixing $A$, so by rigidity of $A\subseteq M(J)$ we have $\Sigma\circ\Theta=\mathrm{id}_{M(J)}$. Similarly,
$(\Theta\circ\Sigma)\circ\jmath\;=\;\Theta\circ\iota\;=\;\jmath,$
so $\Theta\circ\Sigma=\mathrm{id}_{M(F)}$ by rigidity of $(F,\jmath)$. Therefore $\Theta$ is a complete order isomorphism with $\mathrm{ucp}$ inverse $\Sigma$, hence a *-isomorphism $M(J)\xrightarrow{\ \cong\ }M(F)$; in particular,
$(F,\jmath)\ \preccurlyeq\ (J,\iota).$
Thus any rigid cover above $(J,\iota)$ is actually equivalent to it in the order; no strict enlargement exists. Hence $(J,\iota)$ is maximal among rigid covers. ∎
Maximality implies essentiality
Lemma 3.10.
Let $(E,\iota)$ be a maximal rigid $A$-multiplier cover and put $B:=M(E)$. If $\pi:B\to C$ is a complete quotient map such that $\pi|_{A}$ is a complete isometry, then $\pi$ is injective.
Proof.
Let $j:A\hookrightarrow I(A)$ be the Hamana embedding (Theorem 2.1). Since $\pi|_{A}$ is a complete isometry, $\pi(A)$ is completely order isomorphic to $A$ via $\pi|_{A}$. By injectivity of $I(A)$, there exists a ucp map $\Sigma:C\to I(A)$ extending $j\circ(\pi|_{A})^{-1}:\pi(A)\to I(A)$. Compose to obtain $\Theta:=\Sigma\circ\pi:B\to I(A)$, which is ucp and satisfies $\Theta\circ\iota=j$.
Choose any essential ideal $E^{\prime}\lhd I(A)$ (e.g. $E^{\prime}=I(A)$ if unital) and identify $M(E^{\prime})$ with $I(A)$. Then $(E^{\prime},j)$ is a rigid $A$-multiplier cover. The map $\Theta$ shows $(E,\iota)\preccurlyeq(E^{\prime},j)$. If $\pi$ were not injective, then $\Theta$ would not be injective, hence $(E^{\prime},j)$ would strictly dominate $(E,\iota)$, contradicting maximality. Thus $\pi$ is injective. ∎
4. Main theorem: $M(E_{\max})$ realises $I(A)$
Theorem 4.1.
Let $(E_{\max},\iota_{\max})$ be a maximal rigid $A$-multiplier cover. Then
$M(E_{\max})\ \cong\ I(A)$
via the unique complete order isomorphism over $A$.
Proof.
Set $B:=M(E_{\max})$. By Definition 3.1, the inclusion $A\hookrightarrow B$ is rigid. By Lemma 3.10, it is essential in the quotient sense. By Theorem 2.1, $B$ is (canonically) the injective envelope of AA; that is, $B\cong I(A)$ over $A$. ∎
Remark 4.2.
Hamana showed that $I(A)$ need not be a von Neumann algebra; it is in general only monotone complete [7]. Our framework is compatible with this: nothing in the definition of a multiplier cover forces $M(E)$ to be $W^{*}$, and, as noted in Remark 2.3, even if one works inside a $W^{*}$ ambient algebra $B$, one typically gets a quotient $B/\mathrm{Ann}(J)$, which need not be $W^{*}$.
5. The commutative case revisited
We briefly revisit the case $A=C(X)$ with $X$ compact Hausdorff and show how the present isomorphism recovers the classical Gleason cover when expressed via multipliers.
Błaszczyk’s construction proceeds purely topologically:
  1. (a)
    First, we the use Kuratowski–Zorn lemma to get a maximal regular refinement $T^{\ast}$ of the topology of $X$ for which $\mathrm{id}\colon(X,T^{\ast})\to(X,T)$ is irreducible. The resulting $(X,T^{\ast})$ is completely regular and extremally disconnected [1].
  2. (b)
    Then take the Čech–Stone compactification $G(X):=\beta(X,T^{\ast})$. One has that $G(X)$ is extremally disconnected and the induced map $q:G(X)\twoheadrightarrow X$ is irreducible [1].
At this point the $C^{\ast}$-avatar appears: $C(G(X))$ is injective and the inclusion $C(X)\hookrightarrow C(G(X))$ is rigid. Thus $I(C(X))\cong C(G(X))$ by Theorem 2.1.
To realise $I(C(X))$ as a multiplier algebra, work after compactification. Choose any dense open set $U\subseteq G(X)$ (for instance, an increasing union of clopen sets). Then
$M\big(C_{0}(U)\big)\ \cong\ C_{b}(U)\ \cong\ C(\beta U)\ =\ C(G(X)),$
since $\beta U$ identifies with the closure of $U$ inside $G(X)$, which equals $G(X)$ by density. Thus
$I\big(C(X)\big)\ \cong\ M\big(C_{0}(U)\big),$
perfectly mirroring the paradigm maximise first, then compactify. Note that prior to compactification $(X,T^{\ast})$ need not be locally compact, so $C_{0}((X,T^{\ast}))$ is not defined; the multiplier picture belongs naturally to the compact space $G(X)$.
[…]
$M(C_{0}(U))\simeq C(\beta U)=C(G(X))\simeq I(C(X))$.
6. Examples and remarks
Example 6.1.
If $A=C(K)$ with $K$ extremally disconnected or $A=\mathcal{B}(H)$, then $I(A)=A$ and we may take $E=A$ with $\iota=\mathrm{id}$; then $M(E)=E=I(A)$ and the cover is tautologically maximal rigid.
Remark 6.2.
The local multiplier algebra $M_{\mathrm{loc}}(A)$ embeds canonically into $I(A)$ in general, with equality for many classes (e.g. abelian, continuous-trace). The present result shows that $I(A)$ is a multiplier algebra $M(E)$ for a canonical maximal rigid cover $E$.
Remark 6.3.
If one works in a unital ambient algebra $B$ and chooses $J\triangleleft B$, then $M(J)\cong B/\mathrm{Ann}(J)$ (Remark 2.3). Thus $M(J)$ equals $B$ only when $J$ is essential in $B$. In particular, taking $B=\prod_{i}\mathcal{B}(H_{i})$ does not force $M(J)$ to be all of $B$ unless $J$ is essential; and even quotients of $W^{*}$-algebras need not be $W^{*}$ (for example the Calkin algebra is not). This removes the apparent tension with the fact that $I(A)$ is not generally von Neumann [7].
Acknowledgements
We are indebted to Aleksander Błaszczyk for his elegant topological construction [1], which inspired the present point of view.

### v3

@@
 canonically over $A$ by Hamana’s characterisation.
-In the commutative case $A=C(X)$, we recover Błaszczyk’s picture: for a maximal regular refinement $(X,T^{\ast})$ with irreducible identity, $G(X)=\beta(X,T^{\ast})$ and $I(C(X))\cong C(G(X))\cong M(C_{0}(U))$ for any dense open $U\subseteq G(X)$.
+In the commutative case $A=C(X)$, we recover Błaszczyk’s picture: for a maximal regular refinement $(X,T^{\ast})$ with irreducible identity, $G(X)=\beta(X,T^{\ast})$ and $I(C(X))\cong C(G(X))\cong M(C_{0}(U))$ for any dense cozero set $U\subseteq G(X)$ (in particular, for any dense open $F_{\sigma}$ set, e.g.
+an increasing union of clopen sets).
 Thus the paradigm maximise first, then compactify provides a unified conceptual framework on both sides of Gelfand duality.
@@
 Primary 46L05; Secondary 54G05, 46L30
-Institute of Mathematics, Czech Academy of Sciences (RVO: 67985840). Funding received from NCN Sonata-Bis 13 (2023/50/E/ST1/00067) is acknowledged with thanks.
+IM CAS (RVO 67985840).
 1. Introduction
@@
 Gleason’s classical construction [4] of $K$ is intricate; Błaszczyk [1] later gave a strikingly concise route: first maximise the regular topology under an irreducibility constraint, then compactify to obtain $G(X)$.
-Indeed, in Błaszczyk argument, maximality is the driver and extremal disconnectedness is the consequence.
+Indeed, in Błaszczyk’s argument, maximality is the driver and extremal disconnectedness is the consequence.
 Our aim is to transpose this to the noncommutative setting.
@@
 The punchline is that a maximal rigid cover forces $M(E)$ to be a rigid essential extension of $A$, hence identifies canonically with $I(A)$.
+Organisation.
+Section 2 recalls Hamana’s characterisation of $I(A)$ and the multiplier–ideal technology we need (in the correct level of generality).
+Section 3 introduces multiplier covers, defines the forward order by AA‑preserving $\mathrm{ucp}$ maps, and proves that maximality forces essentiality via a lifting argument through $I(A)$.
+It also contains an intrinsic transfinite existence proof of maximal rigid covers using minimal $A^{1}$‑projections and the Choi–Effros product.
+The main identification $M(E_{\max})\cong I(A)$ is established in Section 4, and Section 5 revisits the commutative picture, recovering Błaszczyk.
+We end with remarks on functoriality and local multipliers.
 Dictionary.
@@
 […]
-Organisation.
-Section 2 recalls Hamana’s characterisation of $I(A)$ and the multiplier-ideal technology we need.
-Section 3 introduces multiplier covers, defines the forward order by AA‑preserving $\mathrm{ucp}$ maps, and proves that maximality forces essentiality via a lifting argument through $I(A)$.
-It also contains an intrinsic transfinite existence proof of maximal rigid covers using minimal $A^{1}$‑projections and the Choi-Effros product.
-The main identification $M(E_{\max})\cong I(A)$ is established in Section 4, and Section 5 revisits the commutative picture, recovering Błaszczyk.
-We end with remarks on functoriality and local multipliers.
 2. Preliminaries
@@
 Moreover, $I(A)$ is unique up to a unique complete order isomorphism over $A$. If $A$ is a $C^{\ast}$-algebra, then $I(A)$ is a $C^{\ast}$-algebra and $j$ is a *-monomorphism.
-Equivalently, $I(A)$ is the unique rigid essential extension of $A$: whenever $q:I(A)\to\mathsf{F}$ is a complete quotient map and $q|_{A}$ is a complete isometry, then $q$ is injective.
+Moreover, the Hamana embedding $j:A\hookrightarrow I(A)$ is essential in the following quotient sense: whenever $q:I(A)\to\mathsf{F}$ is a complete quotient map and $q|_{A}$ is a complete isometry, then $q$ is injective.
 2.2. Multiplier algebras and quotienting ideals
@@
 If $Y$ is locally compact Hausdorff, then $M(C_{0}(Y))\cong C_{b}(Y)\cong C(\beta Y)$.
-We shall use the following standard fact.
+We shall use the following standard fact in the form that is valid without additional hypotheses, and record the stronger surjectivity statement under a $\sigma$-unitality assumption.
 Lemma 2.2.
-Let $E$ be a $C^{\ast}$-algebra and $I\lhd M(E)$ a closed two-sided ideal. Put
-[…]
-Then the quotient map $q_{K}:E\to E/K$ induces a surjective *-homomorphism
+Let $E$ be a $C^{\ast}$-algebra and $K\lhd E$ a closed ideal. Then the quotient map $q_{K}:E\to E/K$ extends uniquely to a strictly continuous *-homomorphism
 $\pi_{K}:\ M(E)\longrightarrow M(E/K)$
-whose kernel contains $I$. Conversely, for a closed ideal $K\lhd E$, the strict extension $M(E)\twoheadrightarrow M(E/K)$ has kernel $\{m\in M(E):mE\subseteq K,\ Em\subseteq K\}$.
+whose kernel is $\{m\in M(E):mE\subseteq K,\ Em\subseteq K\}$. The range of $\pi_{K}$ is strictly dense in $M(E/K)$. If $E$ is $\sigma$-unital, then $\pi_{K}$ is surjective.
+Moreover, if $I\lhd M(E)$ is a closed ideal and $K:=\overline{IE+EI}\lhd E$, then $I\subseteq\ker\pi_{K}$.
 Proof.
-See [8, §3.12] and [2]. For $x\in I$ one has $xE,Ex\subseteq K$, hence $x\in\ker\pi_{K}$. The converse kernel description is the usual characterisation of multipliers vanishing modulo $K$. ∎
+See Pedersen [8, §3.12.8–3.12.10] and Busby [2].
+The strict extension and kernel identification are standard; strict density of the range is general, and surjectivity holds when $E$ is $\sigma$-unital.
+The inclusion $I\subseteq\ker\pi_{K}$ follows from $xE,Ex\subseteq K$ for $x\in I$.
+∎
 Remark 2.3.
-If $B$ is unital and $J\lhd B$, the canonical map $B\to M(J)$ given by left/right multiplication is a surjective *-homomorphism with kernel $\mathrm{Ann}(J):=\{b\in B:bJ=Jb=0\}$; hence
-$M(J)\ \cong\ B/\mathrm{Ann}(J).$
-Thus $M(J)=B$ iff $J$ is essential in $B$.
-In particular, even if $B$ is a $W^{*}$-algebra (e.g.
-a product of factors), $M(J)$ need not be $W^{*}$, because $\mathrm{Ann}(J)$ need not be weak-closed.
-This is consistent with Hamana’s examples where $I(A)$ is monotone complete but not von Neumann [7].
+If $B$ is unital and $J\lhd B$, there is a canonical *-homomorphism
+$\lambda_{B,J}:\ B\longrightarrow M(J),\qquad b\mapsto(L_{b},R_{b}),$
+by left/right multiplication; its kernel is $\mathrm{Ann}(J):=\{b\in B:bJ=Jb=0\}$.
+The map $\lambda_{B,J}$ is injective iff $J$ is essential in $B$, but it is not surjective in general (indeed, very rarely; e.g.
+$J=\mathcal{K}(H)$ in its unitisation).
+In particular, even inside a $W^{*}$ ambient algebra there is no automatic von Neumann phenomenon at the level of multipliers.
 3. Multiplier covers and maximality
@@
 An $A^{1}$-projection is an idempotent $A^{1}$-map.
-We endow the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology.
-In what follows, the only continuity we use is:
-$\Phi_{\alpha}\to\Phi\ \ \text{point-ultraweak}\quad\Longrightarrow\quad\Phi_{\alpha}\circ T\ \to\ \Phi\circ T\ \ \text{point-ultraweak}$
-for each bounded linear $T:\mathcal{B}(H)\to\mathcal{B}(H)$; this is automatic because $(\Phi\mapsto\Phi\circ T)$ is continuous when evaluated on fixed vectors.
+We endow the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point–ultraweak topology.
+The only continuity we use is:
+$\Phi_{\alpha}\to\Phi\ \ \text{point–ultraweak}\quad\Longrightarrow\quad\Phi_{\alpha}\circ T\ \to\ \Phi\circ T\ \ \text{point–ultraweak}$
+for each bounded linear $T:\mathcal{B}(H)\to\mathcal{B}(H)$.
 Lemma 3.2.
@@
 $P_{\lambda}\circ P_{\lambda_{0}}\;=\;P_{\lambda}\;=\;P_{\lambda_{0}}\circ P_{\lambda}.$
-Then there exists a ucp map $P_{\infty}:\mathcal{B}(H)\to\mathcal{B}(H)$ with $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$ such that, for every $\lambda_{0}\in\Lambda$,
+Then there exists an $A^{1}$-projection $P_{\infty}:\mathcal{B}(H)\to\mathcal{B}(H)$ such that, for every $\lambda_{0}\in\Lambda$,
 $P_{\infty}\circ P_{\lambda_{0}}\;=\;P_{\infty}\;=\;P_{\lambda_{0}}\circ P_{\infty}.$
@@
 Proof.
-Equip the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology.
-By compactness of the unit ball of completely bounded maps, a subnet of $(P_{\lambda})_{\lambda}$ converges point-ultraweakly to a ucp map $P_{\infty}$; pointwise convergence preserves $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$.
-Fix $\lambda_{0}$ and $x\in\mathcal{B}(H)$.
-Using $P_{\lambda}\circ P_{\lambda_{0}}=P_{\lambda}$ for all $\lambda\geq\lambda_{0}$,
-$P_{\infty}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}(x)\;=\;P_{\infty}(x),$
-so $P_{\infty}\circ P_{\lambda_{0}}=P_{\infty}$.
-A symmetric limit using $P_{\lambda_{0}}\circ P_{\lambda}=P_{\lambda}$ gives $P_{\lambda_{0}}\circ P_{\infty}=P_{\infty}$.
-Hence $P_{\infty}$ is a lower bound.
+By point–ultraweak compactness of the set of ucp maps $\mathcal{B}(H)\to\mathcal{B}(H)$ and closedness of the conditions “unital”, “completely positive”, and “fixes $A^{1}$ pointwise”, there exists a subnet $(P_{\lambda(\alpha)})_{\alpha}$ converging point–ultraweakly to a ucp map $\Phi$ with $\Phi|_{A^{1}}=\mathrm{id}_{A^{1}}$.
+Fix $\lambda_{0}\in\Lambda$ and $x\in\mathcal{B}(H)$.
+For all sufficiently large $\alpha$ we have $\lambda(\alpha)\geq\lambda_{0}$, hence $P_{\lambda(\alpha)}\circ P_{\lambda_{0}}=P_{\lambda(\alpha)}=P_{\lambda_{0}}\circ P_{\lambda(\alpha)}$.
+Passing to the point–ultraweak limit yields
+$\Phi\circ P_{\lambda_{0}}=\Phi=P_{\lambda_{0}}\circ\Phi.$
+Now form the Cesàro averages $A_{N}:=\frac{1}{N}\sum_{k=0}^{N-1}\Phi^{k}$.
+By point–ultraweak compactness, choose a subnet $(A_{N(\beta)})_{\beta}$ converging point–ultraweakly to a ucp map $P_{\infty}$ with $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$.
+Since $(I-\Phi)\circ A_{N}=\frac{1}{N}(I-\Phi^{N})$ and $\|\Phi^{N}\|\leq 1$, we have $(I-\Phi)\circ A_{N}\to 0$ point–ultraweakly, and similarly $A_{N}\circ(I-\Phi)\to 0$.
+Passing to the subnet limit gives
+$\Phi\circ P_{\infty}=P_{\infty}=P_{\infty}\circ\Phi.$
+Hence for every $k\geq 0$, $P_{\infty}\circ\Phi^{k}=P_{\infty}$ and $\Phi^{k}\circ P_{\infty}=P_{\infty}$.
+To see idempotence, note that for each $N$,
+$P_{\infty}\circ A_{N}=\frac{1}{N}\sum_{k=0}^{N-1}P_{\infty}\circ\Phi^{k}=P_{\infty}.$
+Taking $N=N(\beta)$ and passing to the limit gives $P_{\infty}^{2}=P_{\infty}$, so $P_{\infty}$ is an $A^{1}$-projection.
+Finally, from $\Phi\circ P_{\lambda_{0}}=\Phi=P_{\lambda_{0}}\circ\Phi$ we obtain by induction $\Phi^{k}\circ P_{\lambda_{0}}=\Phi^{k}=P_{\lambda_{0}}\circ\Phi^{k}$ for all $k$.
+Thus $A_{N}\circ P_{\lambda_{0}}=A_{N}=P_{\lambda_{0}}\circ A_{N}$ for all $N$, and passing to the subnet limit yields $P_{\infty}\circ P_{\lambda_{0}}=P_{\infty}=P_{\lambda_{0}}\circ P_{\infty}$.
 ∎
@@
 There exists a minimal $A^{1}$–projection $P:\mathcal{B}(H)\to\mathcal{B}(H)$ for the order
-$\Phi\preccurlyeq\Psi\quad\Longleftrightarrow\quad\Phi\circ\Psi=\Phi=\Psi\circ\Phi\quad(\text{equivalently, }\operatorname{Ran}\Phi\subseteq\operatorname{Ran}\Psi).$
-Moreover, any such minimal $P$ is idempotent: $P^{2}=P$.
+$\Phi\preccurlyeq\Psi\quad\Longleftrightarrow\quad\Phi\circ\Psi=\Phi=\Psi\circ\Phi.$
 Proof.
@@
 Lemma 3.4.
-Let $P$ be minimal and let $\Xi$ be any $A^{1}$-projection. Then $P\circ\Xi\circ P=P$.
+Let $P$ be an $A^{1}$-projection on $\mathcal{B}(H)$ and set $R:=P(\mathcal{B}(H))$.
+Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ denote the inclusion.
+If $Q:R\to R$ is a ucp idempotent with $Q|_{A^{1}}=\mathrm{id}$, then
+$\widetilde{Q}\ :=\ \kappa\circ Q\circ P:\mathcal{B}(H)\to\mathcal{B}(H)$
+is an $A^{1}$-projection and satisfies $\widetilde{Q}\preccurlyeq P$.
 Proof.
-Put $F:=P\circ\Xi\circ P$. Then $F$ is an $A^{1}$-projection and $F\preccurlyeq P$, so by minimality $F=P$. ∎
+The map $\widetilde{Q}$ is ucp and fixes $A^{1}$ pointwise since $P$ and $Q$ do and $\kappa$ is unital completely positive. Moreover, using $P\circ\kappa=\mathrm{id}_{R}$ and $Q^{2}=Q$,
+$\widetilde{Q}^{2}=\kappa QP\,\kappa QP=\kappa Q(P\kappa)QP=\kappa QQP=\kappa QP=\widetilde{Q}$.
+Finally,
+$P\circ\widetilde{Q}=P\kappa QP=\kappa QP=\widetilde{Q},\qquad\widetilde{Q}\circ P=\kappa QP^{2}=\kappa QP=\widetilde{Q}$,
+so $\widetilde{Q}\preccurlyeq P$. ∎
 Lemma 3.5.
@@
 Proof.
-Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be the inclusion and set $\widetilde{Q}:=\kappa\circ Q\circ P$. Then $\widetilde{Q}$ is an $A^{1}$-projection, so by Lemma 3.4,
-$P\ =\ P\circ\widetilde{Q}\circ P\ =\ P\circ\kappa\circ Q\circ P\ =\ Q\circ P,$
-using $P\circ\kappa=\mathrm{id}_{R}$ and $P^{2}=P$. For $r=P(x)\in R$ we get $Q(r)=Q(P(x))=P(x)=r$. ∎
+Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be the inclusion and set $\widetilde{Q}:=\kappa\circ Q\circ P$.
+By Lemma 3.4, $\widetilde{Q}$ is an $A^{1}$-projection and $\widetilde{Q}\preccurlyeq P$.
+By minimality of $P$, we have $\widetilde{Q}=P$.
+For $r\in R$ write $r=P(x)$ for some $x\in\mathcal{B}(H)$. Then
+$Q(r)=Q(P(x))=P(x)=r,$
+since $\widetilde{Q}(x)=\kappa(Q(P(x)))=P(x)$. Hence $Q=\mathrm{id}_{R}$. ∎
 Proposition 3.6.
-With $P$ and $R=P(\mathcal{B}(H))$ as above, every ucp map $\theta:R\to R$ with $\theta|_{A^{1}}=\mathrm{id}$ equals $\mathrm{id}_{R}$.
+With $P$ and $R=P(\mathcal{B}(H))$ as above, every $\mathrm{ucp}$ map $\theta:R\to R$ with $\theta|_{A^{1}}=\mathrm{id}$ equals $\mathrm{id}_{R}$.
 Proof.
-Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be inclusion and set $\Xi:=\kappa\circ\theta\circ P$. Then $\Xi$ is an $A^{1}$-projection and by Lemma 3.4,
-$P\ =\ P\circ\Xi\circ P\ =\ P\circ\kappa\circ\theta\circ P\ =\ \theta\circ P.$
-Thus for $r=P(x)\in R$ we have $\theta(r)=\theta(P(x))=P(x)=r$. ∎
+Consider the Cesàro averages $A_{N}:=\frac{1}{N}\sum_{k=0}^{N-1}\theta^{k}$ on $R$.
+By point-ultraweak compactness, a subnet $(A_{N(\beta)})_{\beta}$ converges point-ultraweakly to a ucp map $Q:R\to R$.
+Each $A_{N}$ fixes $A^{1}$ pointwise (since $\theta$ does), hence $Q|_{A^{1}}=\mathrm{id}$.
+From $(I-\theta)\circ A_{N}=\frac{1}{N}(I-\theta^{N})$ and $\|\theta^{N}\|\leq 1$, we have $(I-\theta)\circ A_{N}\to 0$ point-ultraweakly, and similarly $A_{N}\circ(I-\theta)\to 0$.
+Passing to the subnet limit yields
+$Q\circ\theta=Q=\theta\circ Q.$
+In particular, $Q\circ\theta^{k}=Q$ for all $k\geq 0$, so for each $N$,
+$Q\circ A_{N}=\frac{1}{N}\sum_{k=0}^{N-1}Q\circ\theta^{k}=Q.$
+Taking $N=N(\beta)$ and passing to the limit gives $Q^{2}=Q$, so $Q$ is a ucp idempotent on $R$ fixing $A^{1}$. By Lemma 3.5, we get $Q=\mathrm{id}_{R}$.
+Finally, $Q\circ\theta=Q$ with $Q=\mathrm{id}_{R}$ forces $\theta=\mathrm{id}_{R}$. ∎
 Lemma 3.7.
@@
 Proof.
-This is the Choi-Effros theorem [3, Thm. 3.1]. ∎
+This is the Choi–Effros theorem [3, Thm. 3.1]. ∎
 Proposition 3.8.
-Let $P$ be minimal and equip $E:=P(\mathcal{B}(H))$ with the Choi-Effros product. Then:
+Let $P$ be minimal and equip $E:=P(\mathcal{B}(H))$ with the Choi–Effros product. Then:
   1. (a)
@@
 (a) For $a,b\in A^{1}$, $a\odot b=P(ab)=ab$ as $P|_{A^{1}}=\mathrm{id}$. In particular, the restriction to $A$ is a *-homomorphism.
-(b) This is Proposition 3.6 applied to $R=E$.
+(b) This follows from Proposition 3.6 applied to $R=E$.
 (c) Let $\mathcal{S}\subseteq\mathcal{T}$ be operator systems and $u:\mathcal{S}\to E$ ucp.
@@
 Proof.
-Construct $E:=P(\mathcal{B}(H))$ via a minimal $A^{1}$-projection and endow it with the Choi-Effros product as before.
+Construct $E:=P(\mathcal{B}(H))$ via a minimal $A^{1}$-projection and endow it with the Choi–Effros product as before.
 By Proposition 3.8, $A\subseteq E$ is rigid and essential, and $E$ is injective as an operator system.
-Let $J\triangleleft E$ be an essential ideal (take $J=E$ if $E$ is unital).
-Then $M(J)\cong E$ (Remark 2.3), and the inclusion $\iota:A\hookrightarrow M(J)$ yields a rigid $A$-multiplier cover $(J,\iota)$.
+Since $P$ is unital, $E$ is unital, hence $M(E)=E$.
+Set
+J:=E,J\ :=\ E,
+and let $\iota:A\hookrightarrow E=M(E)=M(J)$ be the inclusion (it extends unital to $A^{1}\to E$), which is therefore nondegenerate. Thus $(J,\iota)$ is a rigid $A$-multiplier cover.
 We show $(J,\iota)$ is maximal. Let $(F,\jmath)$ be any rigid $A$-multiplier cover with
 $(J,\iota)\ \preccurlyeq\ (F,\jmath),$
-so there exists an $A$-preserving $\mathrm{ucp}$ map $\Theta:M(J)\to M(F)$ with $\Theta\circ\iota=\jmath$.
-Since $M(J)\cong E$ is injective as an operator system, there is an $A$-preserving $\mathrm{ucp}$ map $\Sigma:M(F)\to M(J)$ with $\Sigma\circ\jmath=\iota$ (Arveson extension along $\jmath$ into $E$).
+so there exists an $A$-preserving $\mathrm{ucp}$ map $\Theta:M(J)=E\to M(F)$ with $\Theta\circ\iota=\jmath$.
+Since $E$ is injective as an operator system, there is an $A$-preserving $\mathrm{ucp}$ map $\Sigma:M(F)\to E$ with $\Sigma\circ\jmath=\iota$.
 Now
 $(\Sigma\circ\Theta)\circ\iota\;=\;\Sigma\circ\jmath\;=\;\iota,$
-hence $\Sigma\circ\Theta$ is a $\mathrm{ucp}$ endomorphism of $M(J)$ fixing $A$, so by rigidity of $A\subseteq M(J)$ we have $\Sigma\circ\Theta=\mathrm{id}_{M(J)}$. Similarly,
+hence $\Sigma\circ\Theta:E\to E$ is $\mathrm{ucp}$ and fixes $A$, so by rigidity of $A\subseteq E$ we have $\Sigma\circ\Theta=\mathrm{id}_{E}$. Similarly,
 $(\Theta\circ\Sigma)\circ\jmath\;=\;\Theta\circ\iota\;=\;\jmath,$
@@
 Let $j:A\hookrightarrow I(A)$ be the Hamana embedding (Theorem 2.1).
-Since $\pi|_{A}$ is a complete isometry, $\pi(A)$ is completely order isomorphic to $A$ via $\pi|_{A}$.
+Since $\pi|_{A}$ is a complete isometry, the operator system $\pi(A)\subseteq C$ is completely order isomorphic to $A$ via $\pi|_{A}$.
 By injectivity of $I(A)$, there exists a ucp map $\Sigma:C\to I(A)$ extending $j\circ(\pi|_{A})^{-1}:\pi(A)\to I(A)$.
-Compose to obtain $\Theta:=\Sigma\circ\pi:B\to I(A)$, which is ucp and satisfies $\Theta\circ\iota=j$.
-Choose any essential ideal $E^{\prime}\lhd I(A)$ (e.g.
-$E^{\prime}=I(A)$ if unital) and identify $M(E^{\prime})$ with $I(A)$.
-Then $(E^{\prime},j)$ is a rigid $A$-multiplier cover.
-The map $\Theta$ shows $(E,\iota)\preccurlyeq(E^{\prime},j)$.
-If $\pi$ were not injective, then $\Theta$ would not be injective, hence $(E^{\prime},j)$ would strictly dominate $(E,\iota)$, contradicting maximality.
-Thus $\pi$ is injective.
-∎
+Set $\Theta:=\Sigma\circ\pi:B\to I(A)$.
+Then $\Theta$ is ucp and $\Theta|_{A}=j$, hence $(E,\iota)\preccurlyeq(I(A),j)$ (viewing $(I(A),j)$ as a rigid $A$-multiplier cover, since $I(A)$ is unital and rigid over $A$).
+By maximality of $(E,\iota)$ among rigid covers, we must also have $(I(A),j)\preccurlyeq(E,\iota)$, so there exists an $A$-preserving ucp map $\Lambda:I(A)\to B$ with $\Lambda\circ j=\iota$.
+Then $(\Lambda\circ\Theta)|_{A}=\mathrm{id}_{A}$, so by rigidity of the cover $(E,\iota)$ we get $\Lambda\circ\Theta=\mathrm{id}_{B}$.
+In particular, $\Theta$ is injective.
+Finally, $\ker\pi\subseteq\ker\Theta=\{0\}$, hence $\pi$ is injective. ∎
 4. Main theorem: $M(E_{\max})$ realises $I(A)$
@@
 Proof.
-Set $B:=M(E_{\max})$.
-By Definition 3.1, the inclusion $A\hookrightarrow B$ is rigid.
-By Lemma 3.10, it is essential in the quotient sense.
-By Theorem 2.1, $B$ is (canonically) the injective envelope of AA; that is, $B\cong I(A)$ over $A$.
-∎
+Set $B:=M(E_{\max})$ and let $j:A\hookrightarrow I(A)$ be the Hamana embedding.
+By injectivity of $I(A)$ as an operator system, the inclusion $j:A\to I(A)$ extends to a ucp map
+$\Theta:B\longrightarrow I(A)\qquad\text{with}\qquad\Theta|_{A}=j.$
+Thus $(E_{\max},\iota_{\max})\preccurlyeq(I(A),j)$ (viewing $(I(A),j)$ as a rigid $A$-multiplier cover).
+Since $(E_{\max},\iota_{\max})$ is maximal among rigid covers, we also have $(I(A),j)\preccurlyeq(E_{\max},\iota_{\max})$. Hence there exists an $A$-preserving ucp map
+$\Lambda:I(A)\longrightarrow B\qquad\text{with}\qquad\Lambda\circ j=\iota_{\max}.$
+Then $(\Lambda\circ\Theta)|_{A}=\mathrm{id}_{A}$, so by rigidity of the cover $(E_{\max},\iota_{\max})$ we obtain
+$\Lambda\circ\Theta=\mathrm{id}_{B}.$
+In particular, $\Theta$ is a unital complete order embedding of $B$ into $I(A)$ which is the identity on $A$.
+Moreover, $\Theta\circ\Lambda:I(A)\to I(A)$ is a ucp idempotent fixing $A$ pointwise, hence $\Theta(I(A))=\Theta(B)$ is an injective operator subsystem of $I(A)$ containing $j(A)$.
+By the minimality property in Theorem 2.1(a), we must have $\Theta(B)=I(A)$.
+Therefore $\Theta$ is surjective, and hence a complete order isomorphism over $A$.
+Since $B$ and $I(A)$ are $C^{\ast}$-algebras, a unital complete order isomorphism is automatically a *-isomorphism. Thus $M(E_{\max})=B\cong I(A)$ canonically over $A$. ∎
 Remark 4.2.
 Hamana showed that $I(A)$ need not be a von Neumann algebra; it is in general only monotone complete [7].
-Our framework is compatible with this: nothing in the definition of a multiplier cover forces $M(E)$ to be $W^{*}$, and, as noted in Remark 2.3, even if one works inside a $W^{*}$ ambient algebra $B$, one typically gets a quotient $B/\mathrm{Ann}(J)$, which need not be $W^{*}$.
+Our framework is compatible with this: nothing in the definition of a multiplier cover forces $M(E)$ to be $W^{*}$, and, as noted in Remark 2.3, even if one works inside a $W^{*}$ ambient algebra $B$, the canonical map $B\to M(J)$ need not be surjective.
 5. The commutative case revisited
@@
   1. (a)
-    First, we the use Kuratowski–Zorn lemma to get a maximal regular refinement $T^{\ast}$ of the topology of $X$ for which $\mathrm{id}\colon(X,T^{\ast})\to(X,T)$ is irreducible.
+    First, we use the Kuratowski–Zorn lemma to get a maximal regular refinement $T^{\ast}$ of the topology of $X$ for which $\mathrm{id}\colon(X,T^{\ast})\to(X,T)$ is irreducible.
 The resulting $(X,T^{\ast})$ is completely regular and extremally disconnected [1].
@@
 At this point the $C^{\ast}$-avatar appears: $C(G(X))$ is injective and the inclusion $C(X)\hookrightarrow C(G(X))$ is rigid. Thus $I(C(X))\cong C(G(X))$ by Theorem 2.1.
-To realise $I(C(X))$ as a multiplier algebra, work after compactification. Choose any dense open set $U\subseteq G(X)$ (for instance, an increasing union of clopen sets). Then
-$M\big(C_{0}(U)\big)\ \cong\ C_{b}(U)\ \cong\ C(\beta U)\ =\ C(G(X)),$
-since $\beta U$ identifies with the closure of $U$ inside $G(X)$, which equals $G(X)$ by density. Thus
+To realise $I(C(X))$ as a multiplier algebra, work after compactification.
+Choose a dense cozero set $U\subseteq G(X)$ (equivalently, a dense open $F_{\sigma}$ set; for instance, an increasing union of clopen sets).
+Then $U$ is locally compact (being open in the compact space $G(X)$), so
+$M\big(C_{0}(U)\big)\ \cong\ C_{b}(U)\ \cong\ C(\beta U).$
+Since $G(X)$ is extremally disconnected compact, it is an $F$-space, and cozero sets are $C^{\ast}$-embedded.
+Consequently every bounded continuous function on $U$ extends uniquely to $G(X)$, so restriction induces an isomorphism $C(G(X))\cong C_{b}(U)$.
+Equivalently, $\beta U\cong G(X)$ (because $U$ is dense and $C^{\ast}$-embedded).
+Thus
+$M\big(C_{0}(U)\big)\ \cong\ C(G(X)).$
+Thus
 $I\big(C(X)\big)\ \cong\ M\big(C_{0}(U)\big),$
@@
 Remark 6.3.
-If one works in a unital ambient algebra $B$ and chooses $J\triangleleft B$, then $M(J)\cong B/\mathrm{Ann}(J)$ (Remark 2.3).
-Thus $M(J)$ equals $B$ only when $J$ is essential in $B$.
-In particular, taking $B=\prod_{i}\mathcal{B}(H_{i})$ does not force $M(J)$ to be all of $B$ unless $J$ is essential; and even quotients of $W^{*}$-algebras need not be $W^{*}$ (for example the Calkin algebra is not).
+If one works in a unital ambient algebra $B$ and chooses $J\triangleleft B$, the canonical map $B\to M(J)$ has kernel $\mathrm{Ann}(J)$ and is injective iff $J$ is essential, but is seldom surjective (for instance, $J=\mathcal{K}(H)$ inside its unitisation).
+Thus even when $B$ is $W^{*}$, $M(J)$ need not be $W^{*}$.
 This removes the apparent tension with the fact that $I(A)$ is not generally von Neumann [7].
-Acknowledgements
-We are indebted to Aleksander Błaszczyk for his elegant topological construction [1], which inspired the present point of view.